Grade 12th passAlgebra

Find the sum:
(5/3) + (3/5)
Find the sum:
(5/3) + (3/5)
Find the sum:
(5/3) + (3/5)

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

Take the LCM of the denominators of the given rational numbers.
LCM of 3 and 5 is 15
Express each of the given rational numbers with the above LCM as the common denominator.
Now,
(5/3)= [(5×5)/ (3×5)] = (25/15)
(3/5)= [(3×3)/ (5×3)] = (9/15)
Then,
= (25/15) + (9/15) ... [∵denominator is same in both the rational numbers]
= (25 + 9)/15
= 34/15

Thanks